The December 31 2018 balance sheet of Justins Golf Shop, INC showed current assets of 1,490, and current liabilities of 1,210. The December 31, 2019 balance sheet showed current assets of 1,675, and current liabilities of 1,290. What was the companys 2019 change in net working capital?
Answer #1

Change in net working capital = Change in current assets - Change in current liabilities

Change in current assets = 1,675 - 1,490 = 185

Change in current liabilities = 1,290 - 1,210 = 80

Change in net working capital = 185 - 80 = 105.

Change in net working capital = 105